# A Simple Latent Diffusion Approach for Panoptic Segmentation and Mask Inpainting

### v1

@@ 4.2 Panoptic Segmentation Results
We ablate the number of inference timesteps in figure 9 and observe that the PQ metric starts plateauing at 20 iterations. Longer inference schedules can further improve the recognition quality (i.e., a reduction in false positives and negatives), while keeping the segmentation quality constant ($\sim 80\%$). We report our best results with 50 iterations.
@@
Figure 9 demonstrates the impact of adding a KL loss to the autoencoding objective in equation 2. It aligns the latents with a standard Gaussian distribution. Unsurprisingly, increasing its loss weight beyond $1e^{-5}$ is not beneficial as it hurts the reconstruction quality. We conclude that weight decay suffices to keep the segmentation latents bounded.
@@
Table 1 verifies the hypothesis that panoptic masks don’t require a powerful VAE. In particular, we finetuned Stable Diffusion’s VAE [59] instead of our shallow autoencoder, as discussed in section 3.1. However, this results in the same final PQ (50.9% vs. 50.8%). Our shallow autoencoder contains 20×20$\times$20 × less parameters, which reduces the training time of stage 2 by 20% (see section 3.2). This observations likely holds for other vision tasks as well.
@@ 5 Conclusion
In summary, this paper presented LDMSeg, a simple yet powerful latent diffusion approach for segmentation and inpainting. It builds upon Stable Diffusion [59] to learn an image-conditioned diffusion process for segmentation in latent space. The experimental results are promising: (i) our framework bypasses specialized object detection modules, such as region proposal networks and bipartite matching, to generate panoptic masks; (ii) our model unlocks sparse panoptic mask completion without finetuning. Hence, we hope that this work will spark further interest into designing general-purpose approaches for dense prediction tasks. Due to the simple and general design of LDMSeg, there is still room for improvement in terms of accuracy and sampling speed. Evident future directions include: training the presented method on larger datasets and extending it to other dense prediction tasks, e.g., depth prediction.
A Supplementary Material
We discuss the implementation details in section B, additional results in section C, limitations in section D, and broader impact in section E.
B Implementation Details
@@
Our best model is trained for 200k iterations on COCO with mixed precision training on 8 $\times$ 40GB NVIDIA A100 GPUs using the Google Cloud. We rely on the pretrained Stable Diffusion [59] weights provided by Hugging Face [24]. We also adopt its settings for the noise scheduler. The code is developed in Pytorch [54] and will be made available as well as our models.
Baseline.
Mask2Former [16] is a specialized segmentation framework that produces excellent results for the panoptic segmentation task. We follow the training recipe from ViTDet [44] and SAM [40] to leverage plain ViT backbones [21] with MAE pretrained weights [30]. Specifically, the model consists of the vision transformer backbone, a shallow neck, and a mask decoder. The latter contains $6$ masked attention decoder layers and $128$ object queries, following [16]. The loss requires Hungarian matching [42] to handle the permutation invariance of the predictions during training. To report the results, we follow its post-processing strategy to combine the classification and mask branches. We adopt the same augmentations as in the main paper, i.e., square resizing and random horizontal flipping. This baseline strikes a good balance between performance, complexity, and training speed. Additionally, we provide results by relying on the backbone and pretrained weights of DINOv2 [53], as we found this to outperform MAE pretrained weights for a ViT-B backbone. We train these models with a batch size of $32$ and a learning rate of $1.5e^{-4}$ for $50$k iterations on 8 $\times$ 16GB V100 GPUs.
@@
Our model produces excellent predictions by simply relying on the $\operatorname*{arg\,max}$ operator. No additional processing is used for the visualizations (see row 3 in figures S2 and S3). In order to report the final PQ metric, however, we eliminate noise by thresholding the predictions at $0.5$ (after applying softmax) and filtering out segments with an area smaller than $512$. These results are shown in the last row of figures S2 and S3. Notice that Mask2Former’s training objective does not impose exclusive pixel assignments, hence it needs additional post-processing steps.
C Additional Results
We show the panoptic segmentation results on COCO val2017 [45] in figure S1 using 50 timesteps. Additionally, we show (class-agnostic) masks in figures S2 and S3. The input images are resized to $3\times 512\times 512$ during training and the diffusion process acts on latents of size $4\times 64\times 64$. To visualize the masks, we assign each segment to a random color. Overall, we conclude that the model is capable of generating high-quality panoptic masks.
@@ C Additional Results
Figure S4 displays the results for different timesteps in the denoising process. Longer sampling benefits the generation of details, such as capturing small objects in the background or an object’s edges. This approach necessitates 10 to 50 iterations to produce high-quality segmentation masks, which is in line with latent diffusion models for images [59]. Furthermore, as the model was forced to distinguish between different instances during training, it’s unlikely that different instances will be grouped during inference. Interestingly, the model iteratively improves the predictions while not reinforcing mistakes during the generative process.
Table 1 provides the inference times for different sampling durations. Unsurprisingly, diffusion models are at least an order of magnitude slower than specialized methods (compare 0.47s vs. 0.02s). However, recent research [65] looks promising to generate high-quality masks in a single step.
@@
Method # Iters Time [s] PQ [%]
Mask2Former 1 0.02 59.0
LDMSeg 10 0.24 50.0
LDMSeg 20 0.47 50.7
LDMSeg 50 1.14 50.8
Table 1: Inference time to generate an image-conditioned mask on a 4090 GPU. The table provides the results for a different amount of denoising steps. We compare to Mask2Former.
D Limitations and Future Work
Undoubtedly, our model has several limitations despite its general design. We discuss two limitations: (i) the model can miss small background objects due to the projection to latent space; (ii) the model is slower during inference than specialized segmentation models due to the adoption of a diffusion prior. In exchange, our method is simple, general, and unlocks out-of-the-box mask inpainting. Finally, increasing the resolution of the latents, enabling open-vocabulary [55] detection, and including additional dense prediction tasks are exciting directions to explore further.
E Broader Impact
The presented approach relies on pretrained weights from Stable Diffusion. Consequently, our model is subject to the same dataset and architectural biases. The user should be aware of these biases and their impact on the generated masks. For instance, these types of (foundation) models have the ability to hallucinate content.

### v2

@@ 4.3 Ablation Studies
  • The impact of the number of inference steps is shown in Figure 9. We observe that the PQ starts plateauing at 20 iterations. Longer inference schedules can further improve the recognition quality (a reduction in false positives and negatives), while keeping the segmentation quality constant ($\sim 80\%$).
@@
  • Figure 9 demonstrates the impact of adding a KL loss to the autoencoding objective in Equation 1. While it aligns the latents with a standard Gaussian distribution, increasing its loss weight beyond $1e^{-5}$ hurts the reconstruction quality. We conclude that weight decay suffices to keep the latents bounded.
@@
  • Table 1 verifies the hypothesis that panoptic masks don’t require a powerful VAE. In particular, we finetuned Stable Diffusion’s VAE [64] instead of our shallow autoencoder (see Section 3.1). However, this results in a similar PQ (50.9% vs. 50.8%). Our shallow autoencoder contains 40×40$\times$40 × less parameters, which reduces the training time of stage 2 up to 20% (see Section 3.2).
@@ 5 Conclusion
We presented LDMSeg, a simple yet powerful latent diffusion approach for panoptic segmentation and mask inpainting. In contrast to prior art, we leverage plain latent diffusion models by building upon Stable Diffusion [64]. In summary, the proposed image-conditioned diffusion process has the following advantages: (i) it bypasses specialized modules, such as region proposals and bipartite matching; (ii) our model unlocks sparse panoptic mask completion without finetuning; (iii) the approach can easily be extended to a multi-task setting by introducing task embeddings. The experiments show that LDMSeg is versatile while also outperforming the majority of prior generalists. Due to its simple and general design, we believe there is still room for improvement in terms of accuracy and sampling speed. Evident future directions include: training LDMSeg on larger datasets and incorporating more dense prediction tasks (e.g., edge detection). Hence, we hope that this work will spark further interest in designing general-purpose approaches for dense prediction tasks.
We discuss the implementation details in Supplement A, additional results in Supplement B, limitations in Supplement C and the broader impact in Supplement D.
Appendix A Implementation Details
@@
Our best model is trained for 100k iterations on COCO with mixed precision training on 8 $\times$ 40GB NVIDIA A100 GPUs using Google Cloud. We rely on the pretrained Stable Diffusion [64] weights provided by Hugging Face [24]. We also adopt its settings for the noise scheduler. The code is developed in Pytorch [59] and will be made available as well as our models.
Multi-Task Setup.
This section provides additional information on the multi-task extension for dense prediction, with minor adaptations. Consider the three fundamental vision tasks: instance segmentation, semantic segmentation and depth prediction. The instance and semantic tasks both utilize the same shallow autoencoder to generate continuous latent codes. Similarly, to compress the depth maps, we rely on the same shallow autoencoder architecture as its segmentation counterpart. We only change the input and output channels to one channel. As COCO does not contain depth annotations, we rely on the predictions from MiDaS [6] to obtain pseudo ground truth. Note that this model predicts relative depth. All tasks use the same set of augmentations and scaling factors, as discussed in the main paper. To enable multi-tasking, we introduce learnable task embedding (786-dimensional) via the cross-attention layers of the UNet. This allows us to query the model for a specific task. Figure S1 visualizes the results for each task by only changing the task embedding. We observe that the model can predict accurate instance, semantic and depth maps for a given image. Finally, given our shallower encoder and task embeddings, a comparison with Marigold [37], a concurrent work on depth estimation, could be insightful.
Figure S1: Multi-Task Setup - Qualitative Results. The figure displays the results for several images in the COCO val set. We can query the model for multiple tasks as it has learned their respective task embeddings.
Mask2Former Baselines.
Mask2Former [15] is a specialized segmentation framework that produces excellent results for panoptic segmentation. We follow the training recipe from ViTDet [46] and SAM [41] to leverage plain ViT backbones [21] with MAE pretrained weights [28]. Specifically, the model consists of the vision transformer backbone, a shallow neck, and a mask decoder. The latter contains $6$ masked attention decoder layers and $128$ object queries, following [15]. The loss requires Hungarian matching [44] to handle the permutation invariance of the predictions during training. To report the results, we follow its post-processing strategy to combine the classification and mask branches. We adopt the same augmentations as in the main paper, i.e., square resizing and random horizontal flipping. This baseline strikes a good balance between performance, complexity, and training speed. Additionally, we provide results by relying on the backbone and pretrained weights of DINOv2 [58], as we found this to outperform MAE pretrained weights for a ViT-B backbone. We train the models with a batch size of $32$ and a learning rate of $1.5e^{-4}$ for $50$k iterations on 8 $\times$ 16GB V100 GPUs.
@@
Our model produces excellent predictions when only relying on the $\operatorname*{arg\,max}$ operator. No additional processing is used for the visualizations (see row 3 in figures S3 and S4). To report the final PQ metric, however, we eliminate noise by thresholding the predictions at $0.5$ (after applying softmax) and filtering out segments with an area smaller than $512$. These results are shown in the last row of figures S3 and S4. Notice that Mask2Former’s training objective does not impose exclusive pixel assignments, hence it needs additional post-processing steps.
Simple Post-Processing.
Panoptic segmentation combines instance and semantic segmentation. After efficiently decoding the latents, we will obtain the panoptic mask by starting from the instances. We subsequently take a majority vote using the predicted semantic mask for each instance. We carry out the following steps and refer to Supplement B for more information on the inference time:
1def postprocess_panoptic(mask_logits, semantic_logits):
2 """
3 Convert predictions to panoptic masks.
4
5 Inputs:
6 mask_logits: np.array of size [N, H, W]
7 semantic_logits: np.array of size [N, H, W]
8 Outputs:
9 panoptic_seg: np.array of size [H, W]
10 segments_to_categories: dict
11 """
12
13 panoptic_seg = np.argmax(mask_logits, axis=0)
14 semantic_seg = np.argmax(class_logits, axis=0)
15
16 segments_ids = {}
17 for segment_id in np.unique(panoptic_seg):
18 instance_mask = panoptic_seg == segment_id
19 if not_confident_or_small(instance_mask):
20 panoptic_seg[instance_mask] = VOID_id
21 continue
22 counts = np.bincount(semantic_seg[instance_mask]])
23 class_id = np.argmax(counts)
24 segments_to_categories[segments_id] = class_id
25
26 return panoptic_seg, segments_ids
Appendix B Additional Results
More Segmentation Results.
We show the panoptic segmentation results with 50 timesteps on COCO val2017 [47] in Figure S2. Additionally, we show (class-agnostic) masks in Figures S3 and S4. The input images are resized to $3\times 512\times 512$ during training and the diffusion process acts on latents of size $4\times 64\times 64$. To visualize the masks, we assign each segment to a random color. Overall, the model is capable of generating high-quality panoptic masks.
@@ Appendix B Additional Results
Number of Denoising Steps.
Figure S5 displays the results for different timesteps during the denoising process. Longer sampling benefits the generation of details, such as capturing small objects in the background or an object’s edges. This approach necessitates 10 - 50 iterations to produce high-quality segmentation masks, which is in line with latent diffusion models for images [64]. Furthermore, as the model was forced to distinguish between different instances during training, it’s unlikely that different instances will be grouped during inference. Interestingly, the model iteratively improves the predictions while not reinforcing mistakes during the generative process.
Inference Time.
Table S2 provides the inference times for different sampling durations. In comparison, Painter requires approximately 0.5 and 0.7 seconds to post-process an image at a resolution of 448 and 560 respectively on our machine. Our post-processing method is significantly faster, taking up only about 0.024 seconds. Importantly, the performance will vary based on hardware and system specifications. Our relatively simple post-processing is explained in Supplement A (final paragraph). Finally, recent research [69] on Consistency Models looks promising to generate high-quality masks in a single step.
Encoding Panoptic Maps.
Table S1 verifies our hypothesis w.r.t. the encoding scheme, as discussed in Sec. 3.1 (main paper). In particular, we test 3 encoding schemes: color (RGB) encoding vs. bit encoding vs. positional encoding:
  • Colors: we generate 256 equidistant colors within the RGB space.
  • Bits: we employ 8 channels to represent integers from $[0,255]$ using bits.
  • Positional: we map integers from $[0,255]$ to an 8-dimensional embedding following [53].
The mIoU and class-agnostic PQ are adopted to measure the reconstruction quality of the autoencoder. We hypothesize that the mapping from color to instance is sub-optimal as this scheme is sensitive to the chosen color palette (89.9 vs. 89.1% PQ). In contrast, bit encoding is a general way to represent discrete panoptic maps, which also outperforms positional encoding (89.9 vs. 88.2% PQ).
Tokenizers and Component Analysis.
Table S3 shows that image tokenizers with more semantically meaningful image features can boost the results. In addition, we show the impact of employing different schedulers and an exponential moving average of the model weights. Note that the results are provided with 50 timesteps during inference. All components further enhance the performance of LDMSeg. To summarize, our best results are obtained with a ViT-B [21] architecture and DINOv2 [58] weights as the image encoder, the DDPM scheduler [31] and an exponential moving average of the model weights during training (weight of $0.999$).
Loss Weights.
@@
Table S1: Encoding. Reconstruction quality for different encoding schemes.
Encoding mIoU PQ [%]
bit encoding 97.3 89.9
color encoding 97.0 89.1
positional encoding 96.7 88.2
Table S2: Inference time. We report the average time to generate a single panoptic mask on COCO with a 4090 GPU. The table provides the results for various denoising steps.
Class-agn. Panoptic Seg. Sem. Seg. Panoptic Seg.
# Iters PQ [%] SQ [%] RQ [%] mIoU [%] PQ [%] SQ [%] RQ [%] Time [s]
1 8.4 76.0 11.1 18.2 8.1 68.9 10.8 0.115
2 35.5 83.9 42.3 21.3 19.8 78.4 24.8 0.160
3 42.4 84.3 50.4 42.1 35.5 79.6 43.8 0.207
4 45.5 84.2 54.0 51.8 39.3 80.3 48.2 0.259
5 47.3 84.1 56.2 55.1 41.3 80.6 50.3 0.320
10 50.2 83.5 60.1 58.6 43.4 80.4 52.6 0.575
15 51.0 83.3 61.2 58.8 43.7 81.3 53.0 0.815
20 51.4 83.2 61.8 59.1 44.1 81.2 53.4 1.071
25 51.7 83.1 62.2 59.6 44.3 81.3 53.7 1.336
30 51.8 83.0 62.4 59.5 44.1 81.0 53.7 1.585
40 52.0 82.9 62.7 59.3 44.3 81.1 53.8 2.062
50 51.9 82.9 62.6 59.9 44.3 81.1 53.8 2.548
60 52.2 82.8 62.8 59.3 44.4 81.2 53.7 3.074
70 52.2 82.7 63.1 59.4 44.3 81.1 53.7 3.564
80 52.2 82.6 63.1 59.3 44.3 80.5 53.8 4.024
90 52.2 82.6 63.1 59.5 44.3 80.5 53.7 4.550
100 52.1 82.7 63.1 59.1 44.3 81.2 53.7 5.030
200 52.1 82.5 63.2 59.1 44.3 80.5 53.7 10.050
Table S3: Component Analysis.
  Setup Image Encoder Scheduler EMA PQ [%]
1 SD VAE [64] DDIM [68] 40.3
2 SD VAE [64] DDIM [68] 40.6
3 ViT-B/14 [21] DDIM [68] 43.7
4 ViT-B/14 [21] DDPM [31] 44.3
Appendix C Limitations and Future Work
Undoubtedly, our model has several limitations despite its general design. We discuss two limitations: (i) the model can miss small background objects due to the projection to latent space; (ii) the model is slower during inference than specialized segmentation models due to the adoption of a diffusion prior. In exchange, our method is simple, general and unlocks out-of-the-box mask inpainting. Moreover, the approach can be extended to a multi-task setting. As we rely on plain diffusion models, new innovations (e.g., architectural, noise scheduler, tokenization, number of inference steps etc.) in image generation are directly applicable to the presented framework. Finally, increasing the dataset’s size, increasing the latents’ resolution, enabling open-vocabulary [60] detection, and including more dense prediction tasks are exciting directions to explore further.
Appendix D Broader Impact
The presented approach relies on pretrained weights from Stable Diffusion [64]. Consequently, our model is subject to the same dataset and architectural biases. The user should be aware of these biases and their impact on the generated masks. For instance, these types of (foundation) models can hallucinate content.